# Boltzmanngasse 5, A 1090 Vienna, Austria

### v2

@@ Appendix B Zeros of $\bar{V}-2\bar{r}\bar{m}$
To see that $\hat{F}(x)$ has at most four real roots we make use of Descartes’ rule of signs, which states that the number of positive real roots of a polynomial is not larger than the number of sign changes in the sequence of polynomial’s non-zero coefficients, while the number of real negative roots is no larger than the number of sign changes after multiplying the coefficients of odd-power terms by $-1$. In our case, we are done if we can show:
@@ Appendix C Regularity of $\bar{H}_{\pm}$
In this appendix we wish to show that the functions $\bar{H}_{\pm}$ of (4.8), i.e.
@@
$\bar{U}=$ $\underbrace{\bigg{(}1+\sum_{i=1}^{N}\frac{{(\bar{\mu}^{i})^{2}}\bar{a}_{i}^{2}}{\bar{r}^{2}+\bar{a}_{i}^{2}}\bigg{)}}_{\breve{U}}\prod_{i=1}^{N}(\bar{r}^{2}+\bar{a}_{i}^{2})=:\breve{U}\bar{F}^{2}\,,$ (C.2)
$N^{2}=$ $\frac{\bar{W}|\bar{V}-2\bar{m}\bar{r}|}{\left(1+\frac{2\bar{m}\bar{r}}{\bar{U}}\sum_{i=1}^{N}\frac{\bar{a}_{i}^{2}(\bar{\mu}^{i})^{2}}{(\bar{r}^{2}+\bar{a}_{i}^{2})\bar{\Xi}_{i}}\right)\prod_{i=1}^{N}(\bar{r}^{2}+\bar{a}_{i}^{2})}$
=\ = $\frac{\breve{U}\bar{W}|\bar{V}-2\bar{m}\bar{r}|}{\breve{U}\bar{F}^{2}+{2\bar{m}\bar{r}}\sum_{i=1}^{N}\frac{\bar{a}_{i}^{2}(\bar{\mu}^{i})^{2}}{(\bar{r}^{2}+\bar{a}_{i}^{2})\bar{\Xi}_{i}}}\,,$ (C.3)
$\bar{W}=$ $1+\sum_{i=1}^{N}\frac{\bar{a}_{i}^{2}(\bar{\mu}^{i})^{2}}{\ell^{2}\bar{\Xi}_{i}}\,,$ (C.4)
$H:=$ $\frac{\breve{U}^{-1}N^{2}}{|\bar{V}-2\bar{m}\bar{r}|}$
=\ = $\frac{1+\sum_{i=1}^{N}\frac{\bar{a}_{i}^{2}(\bar{\mu}^{i})^{2}}{\ell^{2}\bar{\Xi}_{i}}}{\big{(}1+\sum_{i=1}^{N}\frac{{(\bar{\mu}^{i})^{2}}\bar{a}_{i}^{2}}{\bar{r}^{2}+\bar{a}_{i}^{2}}\big{)}\prod_{i=1}^{N}(\bar{r}^{2}+\bar{a}_{i}^{2})+{2\bar{m}\bar{r}}\sum_{i=1}^{N}\frac{\bar{a}_{i}^{2}(\bar{\mu}^{i})^{2}}{(\bar{r}^{2}+\bar{a}_{i}^{2})\bar{\Xi}_{i}}}\,.$ (C.5)
@@
$\lim_{|\bar{\mu}|\to\infty}H\equiv$ $\lim_{|\bar{\mu}|\to\infty}\frac{\breve{U}^{-1}N^{2}}{|\bar{V}-2\bar{m}\bar{r}|}$
=\ = $\frac{\sum_{i=1}^{N}\frac{\bar{a}_{i}^{2}(n^{i})^{2}}{\ell^{2}\bar{\Xi}_{i}}}{\big{(}\sum_{i=1}^{N}\frac{(n^{i})^{2}\bar{a}_{i}^{2}}{\bar{r}^{2}+\bar{a}_{i}^{2}}\big{)}\prod_{i=1}^{N}(\bar{r}^{2}+\bar{a}_{i}^{2})+{2\bar{m}\bar{r}}\sum_{i=1}^{N}\frac{\bar{a}_{i}^{2}(n^{i})^{2}}{(\bar{r}^{2}+\bar{a}_{i}^{2})\bar{\Xi}_{i}}}$
=:absent:=:\ = : $\bar{H}|_{S^{N-1}}\,.$ (C.8)
@@
for some coefficients $\alpha_{i}>0$. Thus both $\vec{X}$ and $\vec{N}$ lie in the positive quadrant of $\mathbb{R}^{N}$. Since all the $\alpha_{i}^{2}a_{i}^{2}$’s are non-vanishing, $\vec{X}$ and $\vec{N}$ are never orthogonal. It follows that the angle between $\vec{X}$ and $\vec{N}$ is in $[0,\pi/2)$. Since $\vec{X}\cdot\vec{N}=|\vec{X}||\vec{N}|\cos(\theta)$ we see that each such scalar product occurring in $\bar{H}|_{S^{N-1}}$ is strictly positive. We conclude that the function $\bar{H}|_{S^{N-1}}(\bar{r},\cdot)$ is bounded both from above and from below for each $\bar{r}$ on the set defined by the inequality
@@
Since $\inf f=-\sup(-f)$, applying what has been established so far to $-\bar{H}$ finishes the proof of the Lemma. $\Box$
Acknowledgements:

### v3

@@ Appendix B Zeros of $\bar{V}-2\bar{r}\bar{m}$
To see that $\hat{F}(x)$ has at most four real roots we make use of Descartes’ rule of signs, which states that the number of positive real roots of a polynomial is not larger than the number of sign changes in the sequence of polynomial’s non-zero coefficients, while the number of real negative roots is no larger than the number of sign changes after multiplying the coefficients of odd-power terms by 1-1. In our case, we are done if we can show:
@@ Appendix C Regularity of $\bar{H}_{\pm}$
In this appendix we show that the functions $\bar{H}_{\pm}$ of (4.8), i.e.
@@
$\bar{U}=\$$\underbrace{\bigg(1+\sum_{i=1}^{N}\frac{{(\bar{\mu}^{i})^{2}}\bar{a}_{i}^{2}}{\bar{r}^{2}+\bar{a}_{i}^{2}}\bigg)}_{\breve{U}}\prod_{i=1}^{N}(\bar{r}^{2}+\bar{a}_{i}^{2})=:\breve{U}\bar{F}^{2}\,$, (C.2)
$N^{2}=\$$\frac{\bar{W}|\bar{V}-2\bar{m}\bar{r}|}{\left(1+\frac{2\bar{m}\bar{r}}{\bar{U}}\sum_{i=1}^{N}\frac{\bar{a}_{i}^{2}(\bar{\mu}^{i})^{2}}{(\bar{r}^{2}+\bar{a}_{i}^{2})\bar{\Xi}_{i}}\right)\prod_{i=1}^{N}(\bar{r}^{2}+\bar{a}_{i}^{2})}$
[…]
$\bar{W}=\$$1+\sum_{i=1}^{N}\frac{\bar{a}_{i}^{2}(\bar{\mu}^{i})^{2}}{\ell^{2}\bar{\Xi}_{i}}\,$, (C.4)
$H:=\$$\frac{\breve{U}^{-1}N^{2}}{|\bar{V}-2\bar{m}\bar{r}|}$
$=\$$\frac{1+\sum_{i=1}^{N}\frac{\bar{a}_{i}^{2}(\bar{\mu}^{i})^{2}}{\ell^{2}\bar{\Xi}_{i}}}{\big(1+\sum_{i=1}^{N}\frac{{(\bar{\mu}^{i})^{2}}\bar{a}_{i}^{2}}{\bar{r}^{2}+\bar{a}_{i}^{2}}\big)\prod_{i=1}^{N}(\bar{r}^{2}+\bar{a}_{i}^{2})+{2\bar{m}\bar{r}}\sum_{i=1}^{N}\frac{\bar{a}_{i}^{2}(\bar{\mu}^{i})^{2}}{(\bar{r}^{2}+\bar{a}_{i}^{2})\bar{\Xi}_{i}}}\,.$ (C.5)
@@
$\lim_{|\bar{\mu}|\to\infty}H\equiv\$$\lim_{|\bar{\mu}|\to\infty}\frac{\breve{U}^{-1}N^{2}}{|\bar{V}-2\bar{m}\bar{r}|}$
[…]
@@
for some coefficients $\alpha_{i}>0$. Thus both $\vec{X}$ and $\vec{N}$ lie in the positive quadrant of $\mathbb{R}^{N}$. Since all the $\alpha_{i}^{2}a_{i}^{2}$’s are non-vanishing, $\vec{X}$ and $\vec{N}$ are never orthogonal. It follows that the angle between $\vec{X}$ and $\vec{N}$ is in $0,\pi/2)$. Since $\vec{X}\cdot\vec{N}=|\vec{X}||\vec{N}|\cos(\theta)$ we see that each such scalar product occurring in $\bar{H}|_{S^{N-1}}$ is strictly positive. We conclude that the function $\bar{H}|_{S^{N-1}}(\bar{r},\cdot)$ is bounded both from above and from below for each $\bar{r}$ on the set defined by the inequality
@@
Since $\inf f=-\sup(-f)$, applying what has been established so far to $\bar{H}$ finishes the proof of the Lemma. $\Box$
Appendix D Analytic continuation
Consider a manifold $\mycal M$} of dimension $d$ equipped with a non-degenerate tensor field $g=g_{\mu\nu}dx^{\mu}dx^{\nu}$ satisfying, say, the vacuum Einstein equations with a cosmological constant,
$\mathrm{Ric}=\lambda g\,$, (D.1)
for some $\lambda\in\mathbb{R}$; the signature is irrelevant for what follows. Suppose that all the metric functions $g_{\mu\nu}$ are real analytic in a coordinate $x^{1}$ near a point $p\in{\mycal M}$. More precisely, let us assume there exists a neighborhood of $p$ with a coordinate system such that $x^{1}(p)=0$ and on which the Taylor series in $x^{1}$, at $x^{1}=0$, of each metric function $g_{\mu\nu}$ converges for all $|x^{1}|<R$ to its value $g_{\mu\nu}(x^{1},\ldots,x^{d})$.
Consider the open disc $D\subset{\mathbb{C}}$ of radius $\rho$, and for $z\in D$ define a complex valued tensor field $\bar{g}_{\mu\nu}(z,x^{2},\ldots,x^{d})$ by the just-mentioned Taylor series. Restricting $\rho$ if necessary we can without loss of generality assume that $|\det(\bar{g})|>0$ on $D$.
We can define the Christoffel symbols, the Riemann tensor, and the Ricci tensor of $\bar{g}$ by the usual formulae of pseudo-Riemannian geometry, with $\partial_{x^{1}}$ replaced by $\partial_{z}$. For real values of $z$ we have $\bar{g}=g$, so that the Ricci tensor of $\bar{g}$, say $\overline{\mathrm{Ric}}$, satisfies there
$\overline{\mathrm{Ric}}=\lambda\bar{g}\,.$
Since $\overline{\mathrm{Ric}}$ is holomorphic in $z$ this remains true for all $z\in D$. For $a\in{\mathbb{C}}$ define
$\psi_{a}(z,x^{2},\ldots,x^{d})=(az,x^{2},\ldots,x^{d})\,.$
Set
$\bar{g}_{a}=\psi_{a}^{*}\bar{g}\,.$
By the usual formulae of tensor calculus, the Ricci tensor $\overline{\mathrm{Ric}}_{a}$ of $\bar{g}_{a}$ satisfies
$\overline{\mathrm{Ric}}_{a}=\psi_{a}^{*}\overline{\mathrm{Ric}}\,$, (D.2)
and we conclude that
$\overline{\mathrm{Ric}}_{a}=\lambda\bar{g}_{a}\,.$ (D.3)
Restricting the coordinate $az$ to real-values, and reverting from $\partial_{z}$ to $\partial_{x^{1}}$, one finds that the Ricci-tensor of the (possibly imaginary-valued) tensor $\bar{g}_{a}$ continues to satisfy (D.3).
Suppose, next, that the metric $\bar{g}_{a}$ above depends analytically on a real-valued parameter, say $b$. A similar argument shows that $\bar{g}_{a}$ can be analytically extended to all values of $b\in{\mathbb{C}}$ which belong to a simply connected and arc-connected Riemann surface on which $|\det g_{a}|>0$, still satisfying (D.3).
One can iterate this argument with further coordinates and parameters in the metric, if applicable.
The task is then to find parameters $a,b\in{\mathbb{C}}$ as above so that the tensor $\bar{g}_{a}$ is real-valued with Lorentzian signature, hence a Lorentzian Einstein metric. In our calculations the coordinates are $t$ and $r$, and the parameter $a$ used is $\sqrt{-1}$ for $t$ and $\sqrt{-1}$ for $r$. The parameter $b$ here corresponds to the $a_{j}$’s and $m$ in the main text.